The return on a stock for the most recent quarter was 2%. Suppose this stock is expected to earn 2% each quarter for the rest of the year.

A. What is the return, stated in terms of an annual percentage rate?

B. What is the return, stated in terms of an effective annual rate?
Answer #1

Return on a stock = 2% in most recent quarter

Expected to ear in each quarter for rest of the year = 2%

a) Annual percentage rate = 2% * 4 = 8%

b) Effective annual interest rate = (1 + (nominal rate / number of compounding periods)) ^ (number of compounding periods) - 1

= ( 1 + 0.08/4) ^ 4 - 1

= 1.08243 - 1

= 8.24%